# ViLLa: Video Reasoning Segmentation with Large Language Model

### v2

@@ 5.3 Main Results
ReferFormer wu2022language 62.9 61.3 64.6 61.1 58.1 64.1
OnlineRefer wu2023onlinerefer 62.9 61.0 64.7 62.4 59.1 65.6
LISA lai2023lisa 63.8 62.1 65.6 62.5 59.2 65.8
PixelLM ren2023pixellm 65.0 63.3 66.8 63.1 60.4 66.0
ViLLa 66.5 64.6 68.6 64.4 61.2 67.7
Table 3: Comparison on VIS datasets (“YTVIS” refers to “Youtube-VIS”).
Method YTVIS-19 YTVIS-21 OVIS
AP AP AP
SeqFormer wu2022seqformer 59.3 51.8 -
Mask2Former cheng2021mask2former 61.6 55.3 24.1
VITA heo2022vita 63.0 57.5 27.7
IDOL wu2022defense 64.3 56.1 42.6
ViLLa 67.1 59.5 46.2
Referring Video Segmentation. Tab. 3 presents the results on the referring video segmentation datasets. Refer-YouTube-VOS seo2020urvos expands Youtube-VOS xu2018youtube using textual descriptions. It consists of 3,975 videos and 27,899 expressions. Refer-DAVIS khoreva2019video extends DAVIS17 pont20172017 , which has 90 videos and more than 1,500 expressions. As shown in the table, our ViLLa demonstrates better performances on these two datasets, surpassing previous referring segmentation models and image-reasoning methods by 3.6 and 1.5 in $\mathcal{J}\&\mathcal{F}$ on Ref-Youtube-VOS dataset, respectively.
Video Instance Segmentation. Tab. 3 presents the results on the video instance segmentation datasets. YouTube-VIS 2019 yang2019video , contains 2.9k videos. The dataset was updated to YouTube-VIS 2021 with longer videos. OVIS dataset is another resource for video instance segmentation, particularly focusing on scenarios with severe occlusions between objects qi2022occluded . It consists of 25 object categories and 607 training videos. Our ViLLa surpasses previous SOTA VIS methods by 2.8, 3.4, and 3.6 points, respectively. The results prove that our model is excelling at modeling temporal relations and segmenting high-quality tracklets.
Table 4: Ablation study on the design choice of vision backbone.
Vision Backbone Segmentation Multiple Choices
AP AR Accuracy
SAM kirillov2023segment 40.5 36.4 46.0
Mask2Former-SwinL cheng2022masked 39.6 35.3 45.7
InternViT  chen2024far 43.5 38.7 52.9
InternVideo2 wang2024internvideo2 44.3 39.2 53.4
Table 5: Ablation study on key component designs in ViLLa.
ID Context Video-Frame VidReasonSeg Refer-VOS
Aggregation Decoder AP Accuracy $\mathcal{J}$$&\mathcal{F}$
1 40.6 49.5 63.7
2 41.5 50.8 64.2
3 44.3 53.4 66.5
Table 6: Ablation study on key designs of Video-Frame Decoder.
ID Multi-scale Video-frame VidReasonSeg Refer-VOS
Tokens Aggregation AP Accuracy $\mathcal{J}$$&\mathcal{F}$
1 41.5 50.8 64.2
2 42.4 50.6 65.0
3 44.3 53.4 66.5
Table 7: Ablation study on the scales.
Architecture VidReasonSeg Refer-VOS
AP Accuracy $\mathcal{J}$$&\mathcal{F}$
Baseline 41.5 50.8 64.2
+ 2 scale frame-video aggregation 43.8 52.5 66.1
+ 3 scale frame-video aggregation 44.3 53.4 66.5
+ 4 scale frame-video aggregation 44.2 53.0 66.4
5.4 Ablation Studies
In this section, we conduct extensive ablation studies to reveal the contribution of each component. Unless otherwise specified, we conduct the experiments on our VideoReasonSeg dataset.
Design Choices of Backbone. We emphasize that different vision backbones can be applicable in our framework. In Tab. 5, we show that Internvideo2 performs the best, potentially because it is the state-of-the-art video foundation model, demonstrating its capacity in multimodal video understanding. Further, with the Mask2Former backbone, our framework still achieves a decent performance on the video reasoning segmentation dataset. This reveals that the design choice of the vision backbone is flexible and not limited to the video foundation model.
Key Components Design. Tab. 5 demonstrates the effect of our component designs based on the vision backbone and transformer decoder in Mask2Former. By adopting our algorithm, we significantly improve the baseline model by strengthening the video-frame relationship and fusing context information into text embedding. Tested on VideoReasonSeg, our context aggregation improves the accuracy of answering multiple choices by 1.3 points, which indicates that adding context information to the text embedding and keeping the textual input contextualized is beneficial for video understanding. Also, the design of video-frame decoder improves the segmentation performance by 2.7 AP on segmentation, showing that improving the video modeling by applying multi-scale tokens and strengthening their relations is crucial.
Video-Frame Decoder. Tab. 7 shows the influence of utilizing multi-scale tokens and video-frame aggregation in video-frame decoder. With multi-scale tokens only, the baseline model achieves a 0.9 AP increase, while applying the aggregation brings about 1.9 AP. This demonstrates that using multi-scale tokens and associating them are beneficial for video segmentation tasks. In Tab. 7, we explore the effects of increasing the number of scales in the decoder. Initially adding layers will contribute to notable gains, but the gain would diminish with more scales of decoder layers.
Figure 3: Comparison between ViLLa, LISA, and PixelLM.
5.5 Qualitative Results
In Fig. 3, we provide a visual comparison with existing related works. We show that our ViLLa is capable of segmenting multiple instances (helmets of riders) while tracking them. On the other hand, LISA and PixelLM, though aggregated with the tracker, can not either segment multiple instances or track all instances properly (as shown in red circles, where they both track the wrong helmet). These qualitative results demonstrate the reasoning and segmentation capacity of our ViLLa.
6 Conclusion
In this study, we propose a new task, video reasoning segmentation. Also, we have introduced a benchmark VideoReasonSeg, which comprises approximately 2k samples. Finally, we introduce ViLLa, an effective and efficient model that aims to provide some early explorations towards this task. Benefiting from our exquisite designs, ViLLa demonstrates convincing capabilities in producing temporally consistent high-quality masks for a variety of video segmentation tasks.
Limitations and broader impacts. Even though our model performs well in reasoning segmentation with text instructions, the user input is still limited to text, rather than a variety, such as scribbles. Also, our model can not support multi-round conversations. However, We hope our work can provide new insights into the future direction of combining LLMs and vision tasks so as to build an omnipotent perception model.
Acknowledgement. This work is partially supported by the National Natural Science Foundation of China (No. 62201484), National Key R&D Program of China (No. 2022ZD0160100), HKU Startup Fund, Shanghai Artificial Intelligence Laboratory, and HKU Seed Fund for Basic Research.
Appendix
This appendix provides more details about the proposed ViLLa, and our proposed benchmark, VideoReasonSeg. The first part includes data details while the second part contains the data generation pipeline of our VideoReasonSeg dataset, and in the final part we provide extra ablation experiment.
Appendix A Implementation Details
@@
Appendix B VideoReasonSeg Details
Data Generation Pipeline.
As for the question and answer pair, we use GPT-4V and Qwen-VL to construct our dataset. We utilize videos with pre-existing video mask annotations. The video frames, the category names contained in the video, and their related mask annotation are contained in the prompts feed to the GPT-4V. An example of the prompts is shown in Fig. 4. Using carefully crafted prompts, GPT-4V autonomously selects instances to construct question-answer pairs relevant to the video. As illustrated in Fig. 5, we demonstrate two types of question, both question and multiple choice, from a given video. In this example, we show that our data tests the capability of models of reasoning based on common world-knowledge, and relate ‘vehicle carrying passengers’ to the white bus on the roadside. In addition, the multiple choice expects the model to distinguish the type of vehicle from other plausible answers, such as ‘taxi’ and ‘bike’. All together these questions are tests of the reasoning capacities of models on both pixel-level and video-level.
Although GPT-4V can efficiently understand the content of the video frame, there are still failure cases in the generated data. One major problem is that questions can be too objective and hard to evaluate. For example, the question “How would you rate the overall difficulty and impressiveness of the skateboarding you observed?” is very objective and the answers can be various for different people. And this requires further prompts and filtering during the data generation process.
Visualizations of ViLLa. Extra visualizations are presented in Fig. 6, which show a case that successfully segments a brown cow moving forward.
Figure 5: GPT-4V data generation samples. The part shows further samples of the generated questions and the multiple choices. The two types of questions can help we better evaluate the model’s performance in video reasoning from both pixel-level and video-level.
Figure 6: Visualizations of ViLLa.
Appendix C Ablation Studies
Aggregation Strategy. Tab. 9 shows the results of different aggregation strategies in the video-frame decoder. We compare our aggregation between video-level segmentation embeddings with the feature fusion adopted in PixelLM [36]. As shown in the table, our strategy improves the performance on referring VOS dataset, demonstrating the effectiveness of the video-frame aggregation strategy.

### v3

@@ Appendix A Discussions
ViLLa 51.9 44.0 47.0 43.9 50.1
ViLLa 55.4 49.9 49.4 46.5 52.3
@@
Appendix B Implementation Details
Appendix C Additional Ablation Experiments
Sampling Strategy. In the key segment extractor, we take the average from the top-K responses to obtain the starting and ending frames of the key segments, where we denote as $V_{key}$, comprising $T_{key}$ frames. Based on the key segments, we also sample $T_{ref}$ using an adaptive global sampling strategy. ‘Global’ indicates sampling $T_{ref}$ frames from the whole video apart from the key segments we extract, and ‘neighbor’ denotes sampling frames from the neighboring frames (both precedent and antecedent) of the key segments. Our ‘adaptive’ sampling strategy, on the other hand, combines both ‘global’ and ‘local’ sampling strategies, which samples $T_{ref}/3$ from the whole video, and $2/3T_{ref}$ from the neighboring frames. As shown in Tab. 10, adaptive sampling slightly outperforms both global and neighbor sampling strategies. As the number $T_{ref}$ increases, the performance gradually improves.
Aggregation Strategy. Tab. 11 shows the results of different aggregation strategies in the segment synchronization decoder. We compare our aggregation between video-level segmentation embeddings with the feature fusion adopted in PixelLM [46]. As shown in the table, our strategy improves the performance on referring VOS dataset, demonstrating the effectiveness of the video-frame aggregation strategy.
Video Instance Segmentation. Tab. 12 presents the results on the video instance segmentation datasets. YouTube-VIS 2019 [70], contains 2.9k videos. The dataset was updated to YouTube-VIS 2021 with longer videos. OVIS dataset is another resource for video instance segmentation, particularly focusing on scenarios with severe occlusions between objects [43]. It consists of 25 object categories and 607 training videos. Our ViLLa surpasses previous SOTA VIS methods by 3.3, 3.8, and 3.9 points, respectively. The results prove that our model is excelling at modeling temporal relations and segmenting high-quality tracklets.
Table 10: Ablation study on different sampling strategies.
      Strategy       $T_{ref}$       ReasonVideoSeg
      $\mathcal{J}$$&\mathcal{F}$       Accuracy
      Global       0       54.0       47.8
      6       54.3       48.3
      12       54.5       48.7
      Neighbor       0       54.0       47.8
      6       54.4       48.5
      12       54.7       49.2
      Adaptive       0       54.0       47.8
      6       54.8       49.0
      12       55.4       49.9
@@ Appendix C Additional Ablation Experiments
Table 12: Comparison on multiple VIS datasets.
Method YTVIS-19 YTVIS-21 OVIS
AP AP AP
SeqFormer [66] 59.3 51.8 -
Mask2Former [7] 61.6 55.3 24.1
VITA [21] 63.0 57.5 27.7
IDOL [67] 64.3 56.1 42.6
ViLLa 67.6 59.9 46.5
Appendix D VideoReasonSeg Details
As for the question and answer pair, we use GPT-4V to construct our dataset. We utilize videos with pre-existing video mask annotations. The video frames, the category names contained in the video, and their related mask annotation are contained in the prompts fed to the GPT-4V. An example of the prompts is shown in Fig. 4. Using carefully crafted prompts, GPT-4V autonomously selects instances to construct question-answer pairs relevant to the video. As illustrated in Fig. 5, we demonstrate two types of questions, both question and multiple choice, from a given video. In this example, we show that our data tests the capability of models of reasoning based on common world-knowledge, and relate ‘vehicle carrying passengers’ to the white bus on the roadside. In addition, the multiple choice expects the model to distinguish the type of vehicle from other plausible answers, such as ‘taxi’ and ‘bike’. All together these questions are tests of the reasoning capacities of models on both pixel-level and video-level.
Although GPT-4V can efficiently understand the content of the video frame, there are still failure cases in the generated data. One major problem is that questions can be too objective and hard to evaluate. For example, the question “How would you rate the overall difficulty and impressiveness of the skateboarding you observed?” is very objective, and the answers can vary for different people. This requires further prompts and filtering during data generation process.
Video visualizations of additional data of the VideoReasonSeg are further presented in Fig. 6, which shows varied cases that include: a) discrimination from multiple instances; b) multiple instances with fast movement; c) open-world knowledge reasoning.
Appendix E Failure Case Analysis
Even though our ViLLa shows impressive results in video reasoning segmentation, there is still room for improvement. As shown in Fig. 7, ViLLa incorrectly segments the bystander who is observing the two-person talking. We hypothesize that this error arises from the inability of the MLLM to temporally localize the “talking” action, which occurs exclusively in the final three frames (there is even an occlusion in the last frame). Consequently, ViLLa erroneously associates the individual co-occurring with the motor-riding man in the initial frames and persistently tracks this subject throughout the whole video.
Figure 5: GPT-4V data generation samples. The part shows further samples of the generated questions and the multiple choices. The two types of questions can help us better evaluate the model’s performance in video reasoning at both pixel-level and video-level.
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
Figure 6: Data samples of our proposed VideoReasonSeg.
Figure 7: Failure case. ViLLa incorrectly segments the bystander who is observing the two-person talking.